What pH is required to remove all but 0.3 mg/L of the Aluminum ion?

Prairie View, TX groundwater contains 2.2 mg/L of Aluminum (Al3). What pH is required to removed all but 0.3 mg/L of this ion at 25°C. (use pKs 7. 32.9)
Answer #1

Given the pKsp of the reaction as 32.9

pKsp = -log10 (Ksp)

Thus Ksp = 10-32.9

Al(OH)3 (s)\rightleftharpoons Al+3 + 3OH- Ksp = 10-32.9

[We assigned Ksp to forward reaction since Al(OH)3 (s) formation is more favourable than dissociation in to Al+3, OH- ]

Ksp = ([Al+3]* [OH-]3)/Al(OH)3 (s)

Since [Al(OH)3 (s)] is a precipitate and its concentration is taken as 1

Thus  Ksp = [Al+3]* [OH-]3

10-32.9 = [Al+3]* [OH-]3 -----------(1)

Given the ground water contains 2.2 mg/l of Al+3

and we want to remove 1.9 mg/l (2.2-0.3) of Al+3

Al+3 = 1.9 mg/l = (1.9 * 10-3 / 26.981) mol/l = 7.042*10-5 mol/l [since molecular weight of Al+3 = 26.981 g/mol]

Substitute Al+3 =  7.042*10-5 mol/l in equation (1) (since this amount of aluminium ion has to react with OH- to produce Al(OH)3 precipitate)

10-32.9 =  7.042*10-5 mol/l * [OH-]3

[OH-]3 = 10-32.9/ 7.042*10-5

[OH-] = (10-32.9/ 7.042*10-5)(1/3)

[OH-] = 2.6147 * 10-10 mol/l ------------(2)

As we know for water Reaction Kw = 10-14 = [H+] * [OH-] ---------------(3)

Substituting (2) in (3)

we get [H+] = 3.8244 * 10-5 mol/l

pH = -log10[H+] = -log10[3.8244 * 10-5] = 4.417

Thus the required pH of ground water to remove 1.9 mg/l of Al+3 is 4.417